A man pushes a box of 40 kg up incline plane of 15°, if the man applies a horizontal force 200N and the box moves up the plane a distance of 20 m at a constant velocity and the coefficient of friction is 0.1, find
1)The work done by the man
2)And also find the work done against friction
1
Expert's answer
2019-02-28T12:32:08-0500

1) The work done by the man

W=Fd cos⁡〖θ=〗 200 N×20 m×cos⁡〖15°=3864 J〗

2) The work done by the friction

W_f=-W+mgh=-W+mgd sin⁡θ

=-3864 J+40 kg×9.8 N/kg×20 m×sin⁡〖15°=-1835 J〗